# The Monte Carlo Computational Summit – October 25 & 26, 2023 – Notre Dame, Indiana, USA

### v1

@@ 3.8 Experience Adapting OpenMC to CUDA
In parallel to efforts at Argonne National Laboratory to adapt OpenMC for GPU computation, a smaller thrust at the Massachusetts Institute of Technology focused on a port using the more widely adopted CUDA toolchain (Ridley and Forget, 2021). The software development sought to minimize the source code change between the CPU-based code and the GPU-based code, potentially incurring a performance cost. About 50,000 lines of C++ source code comprise the core operations which must be accelerated by GPU.
@@
After applying the cache-less microscopic cross section trick mentioned in the prior OpenMC section and by Tramm et al. (2022), along with a variety of other optimizations such as trading occupancy for reduced thread divergence, the CUDA-based code achieved a performance about a factor of two worse than the OpenMP code. Given the widespread use of Monte Carlo transport codes, our recommendation for Monte Carlo particle transport code developers going forward is to use flat data structures without polymorphism. Since optimal performance is important here, the extra programmer effort required to avoid software abstraction pays off in the long run.
@@ 3.9 Monte Carlo Research at NNL
MC21 is the Naval Nuclear Laboratory’s primary Monte Carlo code. It is used internally for reactor physics, criticality safety, and shielding analysis and is written primarily as a production design tool (Griesheimer et al., 2015). MC21 has support for:
@@
GPU support in MC21 is only available in a developmental fork and parallelism schemes are governed by machines available at NNL. MC21 currently uses both OpenMP and MPI parallelism, primarily targeting Intel Xeon-type machines. Using this tact, simulations have been performed at the peta-scale. When approaching GPU porting, the MC21 development team has focused on minimizing code touched by a port and maintaining compatibility with in-line feedback, searches, and tally systems. The team started by selectively porting only bottle-necking kernels beginning with macroscopic cross section lookups for tracking (Klinvex et al., 2021). Tracking and tallying is done on the CPU, then when a macroscopic cross section is needed that particle is banked and tracking is paused. This continues until a sufficient number of cross section lookup calls is batched then they are executed all at once on the GPU. While mini-app results were promising, when implemented in the full code only a 10% speedup in the most-lookup-bound case was found with a 30% slowdown for more representative simulations.
@@ 4 Standardization of Test Problems
By the end of the discussion, the majority of summit members determined a time-dependent continuous-energy form of the C5G7 challenge problem (Hou et al., 2017) to be a suitable candidate for a standardized test problem. A unified figure of merit was identified: average reaction tallies across each pin region of the problem; the problem would be modified by developers at Oak Ridge National Laboratory to allow for this figure of merit to be reported. Additional features, such as multi-group and implicit capture, are also to be added as toggle-able options to the base problem; however, the base problem was chosen as the standard for all Monte Carlo codes for the reporting of comparable figures of merit.
@@ 6 Acknowledgments
We would like to thank the attendees, presenters, and organizers including co-chairs Alex Long and Steven Hamilton. We would also like to thank the University of Notre Dame and ND Energy for hosting the Monte Carlo computational summit.
Joanna Piper Morgan currently holds a co-op position at Advanced Micro Devices (AMD) in addition to her position with CEMeNT at Oregon State University and has previously worked at Los Alamos National Laboratory on MCATK. All other authors declare no conflicts of interest.

### v2

@@ 3.8 Experience Adapting OpenMC to CUDA
In parallel to efforts at Argonne National Laboratory to adapt OpenMC for GPU computation, a smaller thrust at the Massachusetts Institute of Technology focused on a port using the more widely adopted CUDA toolchain [39]. The software development sought to minimize the source code change between the CPU-based code and the GPU-based code, potentially incurring a performance cost. About 50,000 lines of C++ source code comprise the core operations which must be accelerated by GPU.
@@
After applying the cache-less microscopic cross section trick mentioned in the prior OpenMC section and by [38], along with a variety of other optimizations such as trading occupancy for reduced thread divergence, the CUDA-based code achieved a performance about a factor of two worse than the OpenMP code. Given the widespread use of Monte Carlo transport codes, our recommendation for Monte Carlo particle transport code developers going forward is to use flat data structures without polymorphism. Since optimal performance is important here, the extra programmer effort required to avoid software abstraction pays off in the long run.
@@ 3.9 Monte Carlo Research at NNL
MC21 is the Naval Nuclear Laboratory’s primary Monte Carlo code. It is used internally for reactor physics, criticality safety, and shielding analysis and is written primarily as a production design tool [40]. MC21 has support for:
@@
GPU support in MC21 is only available in a developmental fork and parallelism schemes are governed by machines available at NNL. MC21 currently uses both OpenMP and MPI parallelism, primarily targeting Intel Xeon-type machines. Using this tact, simulations have been performed at the peta-scale. When approaching GPU porting, the MC21 development team has focused on minimizing code touched by a port and maintaining compatibility with in-line feedback, searches, and tally systems. The team started by selectively porting only bottle-necking kernels beginning with macroscopic cross section lookups for tracking [41]. Tracking and tallying is done on the CPU, then when a macroscopic cross section is needed that particle is banked and tracking is paused. This continues until a sufficient number of cross section lookup calls is batched then they are executed all at once on the GPU. While mini-app results were promising, when implemented in the full code only a 10% speedup in the most-lookup-bound case was found with a 30% slowdown for more representative simulations.
@@ 4 Standardization of Test Problems
By the end of the discussion, the majority of summit members determined a time-dependent continuous-energy form of the C5G7 challenge problem [29] to be a suitable candidate for a standardized test problem. A unified figure of merit was identified: average reaction tallies across each pin region of the problem; the problem would be modified by developers at Oak Ridge National Laboratory to allow for this figure of merit to be reported. Additional features, such as multi-group and implicit capture, are also to be added as toggle-able options to the base problem; however, the base problem was chosen as the standard for all Monte Carlo codes for the reporting of comparable figures of merit.
@@ 6 Acknowledgments
We would like to thank the attendees, presenters, and organizers including co-chairs Alex Long and Steven Hamilton. We would also like to thank the University of Notre Dame and ND Energy for hosting the Monte Carlo computational summit. Joanna Piper Morgan has previously had a co-op position at Advanced Micro Devices (AMD) and previosuly interned at Los Alamos National Laboratory where she worked on MCATK. All other authors declare no conflicts of interest.